OpenStudy (anonymous):

Prove that the series \(\Large{exp(x) = \sum_{n=0}^{\infty} \frac{x^{n}}{n!}}\) does not uniform converge on \(\mathbb{R}\), although the limit function (convergence radius \(R=+\infty\)) is continous on \(\mathbb{R}\).

OpenStudy (kinggeorge):

From what it seems like to me, we need to show that the sequence of partial sums is not uniformly convergent. How we would show that, I'm not entirely sure yet.

OpenStudy (anonymous):

hmm.. maybe Mr Elias has idea for it..

OpenStudy (anonymous):

Let \[s_n(x) =\sum_1^n \frac {x^i}{i!} \] Suppose the the convergence is uniform. That means that the is N such that for n >= N such that \[ |s_n(x) -e^x| < 1, \text { for all } x\\ \text { in particular } |s_N(x) -e^x| < 1, \text { for all } x\\ \] This is not possible since \[ \lim_{x \to \infty} S_N(x)= + \infty \]

OpenStudy (anonymous):

I still have to make a better proof.

OpenStudy (anonymous):

Here how to make it works. Let N odd, then \[ \lim_{n\to - \infty} S_N(x)= -\infty \\ | -\infty -0 | <1 \] Which is a contradiction

OpenStudy (anonymous):

Mr Elias it looks good but if you want to make better proof i will be happier, because my mentor is very mean about giving points.. maybe George can say something additional

OpenStudy (anonymous):

It is fixed now, See my previous post.

OpenStudy (anonymous):

ok thank you very much Mr Elias,

OpenStudy (anonymous):

yw